OpenStudy (anonymous):

A data set consists of the following data points:(2,12), (3,7), (5,4)The slope of the best fit line is -2.5. Find the y-intercept of this line?

OpenStudy (anonymous):

can someone figure this out for me?

OpenStudy (amistre64):

what is your average x value? and your average y value?

OpenStudy (anonymous):

I have like no idea...

OpenStudy (amistre64):

\[(y-\bar y)=m(x-\bar x)\] \[y=-2.5(x-\bar x)+\bar y\] \[y=-2.5(0-\bar x)+\bar y\] \[y=2.5\bar x+\bar y=yintercept\]

OpenStudy (amistre64):

if you do not know how to find the average of a set of numbers, then doing linear regression is not your cup of tea yet

OpenStudy (anonymous):

it isn't its a online class that im totally lost in..

OpenStudy (amistre64):

how do you find the average of a list of numbers? you have to have some idea about it ....

OpenStudy (anonymous):

don't you just add them all together then divide them by the amount of numbers their are like (1+1+1+1)/4?

OpenStudy (amistre64):

perfect :) add all the x values, and divide add all the y values, and divide the anchoring point for a set of points is the average(x,y) values

OpenStudy (amistre64):

its the slope of the linear regression thats a pain by hand, but they were nice enough to give that to you

OpenStudy (anonymous):

so for the x values It would be ( 2+3+5)/3=3.33 y-values(12+7+4)/3=7.667?

OpenStudy (amistre64):

correct; and i already worked the point slope form of a line down to its y intercept y int = 2.5(X) + (Y) = 2.5(3.333) + (7.667)

OpenStudy (amistre64):

if you dont approximate it, but use 10/3 and 23/3 it comes out as an integer

OpenStudy (anonymous):

ohh I just added it up and got 15.9995? what did I do wrong?

OpenStudy (amistre64):

thats fine sine you approximated the X and Y averages. you introduced an error by aprroximating instead of using exact values. it turns out to be 16, which is fine when you rnd up

OpenStudy (anonymous):

what!!! I kind of got it lol thanks for helping me out.!!

OpenStudy (amistre64):

:) good luck